Find the slope-intercept equation for the line with the indicated slope and y-intercept. Slope 3/4; y-intercept (0, 35)
Answer by alicealc(293)

slope (m) = 3/4
intercept (x1, y1) = (0, 35)

this formula can be used to find the equation:
y - y1 = m * (x - x1)
y - 35 = 3/4 * (x - 0)
y - 35 = 3/4 * x
multiply both sides by 4:
4y - 140 = 3x
3x - 4y + 140 = 0